# The free energy of Dirac’s vacuum in purely magnetic fields

### v1

@@ Abstract.
The QED vacuum is a non-linear polarisable medium rather than an empty space. In this work, we present the first rigorous derivation of the one-loop effective magnetic Lagrangian at positive temperature, a non-linear functional describing the free energy of quantum vacuum in a classical magnetic field.
We start by properly defining the free energy functional using the Pauli-Villars regularisation technique in order to remove the worst ultraviolet divergence which represents a well known issue of the theory. Then, we study the limit of slowly varying classical magnetic fields. In this regime, we prove the convergence of this functional to the Euler-Heisenberg formula with thermal corrections, by recovering the effective Lagrangian first derived by Dittrich [16] in $1979$.
@@ 1. Introduction
The non-linear polarisable nature of the vacuum in quantum field theory has been known for a long time [14, 22, 55]. In this framework, the classical action is replaced by an effective action taking the quantum corrections into account while guaranteeing the validity of the principle of least action. In quantum electrodynamics, a standard way to get the effective action is by integrating out fermions in the path integral (see for instance [47, Ch. 33]). In this manner, we obtain a functional of a classical electromagnetic field treated as an external one. In [28], the authors derive in a rigorous way the effective Lagrangian action for time-independent fields in the Coulomb gauge,
$\mathcal{L}\left(\boldsymbol{A}\right)=-\mathcal{F}_{\mathrm{vac}}\left(e\boldsymbol{A}\right)+e\int_{\mathbb{R}^{3}}\left(j_{\mathrm{ext}}\cdot A\left(x\right)-\rho_{\mathrm{ext}}\left(x\right)V\left(x\right)\right)\mathop{}\mathopen{}\mathrm{d}x+\frac{1}{8\pi}\int_{\mathbb{R}^{3}}\left(|E\left(x\right)|^{2}-|B\left(x\right)|^{2}\right)\mathop{}\mathopen{}\mathrm{d}x,$
@@
where $f_{\mathrm{vac}}\left(eE,eB\right)$ is the energy per unit volume for a constant electromagnetic field $\boldsymbol{F}=\left(E,B\right)$. The function $f_{\mathrm{vac}}$ was first computed by Euler and Heisenberg [37] in $1936$. Some alternative derivations were then presented by Weisskopf [55] and Schwinger [48], who introduced a largely renowned method in the physics literature known as proper time method. Since the integrand in the Euler-Heisenberg formula [31, Formula 4] can have poles on the real line, a proper definition needs to rotate the integration path into the complex plane by replacing the integral variable $s$ by $s+i\eta$ and taking the limit $\eta\xrightarrow{}0$ [48]. In this manner, the function $f_{\mathrm{vac}}$ gains an exponentially small non-zero imaginary part which was interpreted by Schwinger as the electron-positron pair production rate and explains the instability of the vacuum [17][32, Paragraph 7.3]. Nevertheless, in the case of a purely magnetic field, the Euler-Heisenberg formula is absolutely convergent and real, as expected. For this reason, this situation has been deeply studied in the physics literature [7, 51, 52, 41] and the function $f_{\mathrm{vac}}$ can be simplified as follows:
@@
In the non-perturbative physical examples above, the fields exist in a thermal bath or in some non-equilibrium background which is very different from the vacuum. For this reason, the study of finite temperature effects in quantum field theory attracted the interest of many physicists in the past decades. In particular, several authors [38, 54, 18] considered what happens when a system of elementary particles described by a quantum field theory is heated. They found that symmetries which are spontaneously broken at zero temperature (such as those of the weak interactions) may be restored at sufficiently high temperatures, and calculated the critical temperature at which such a restoration takes place. To do this kind of calculation, one needs to know the Feynman rules for a field theory at finite temperature. For a non-gauge theory, these rules can be derived using well known methods [24]. However, for a gauge theory, a more powerful technique is needed to cope with several new problems that arise [4]. In any case, the Feynman rules at finite temperature $T$ consists in the zero temperature rules with the following formal replacements:
@@
It is now clear that the calculation of the effective action at positive temperature has been an interesting physical question for some decades, particularly because the imaginary part of the effective action is expected to give the vacuum decay rate and one would like to understand the influence of the temperature on this rate. There have been several attempts to generalise Schwinger’s proper time method to the finite temperature case in order to determine the temperature dependent effective actions, leading to conflicting results in the physics literature. The subtleties behind these results have to do with the imposition of the anti-periodicity condition required at positive temperature as well as on the finite temperature formalism used to carry out the calculation. We refer the interested reader to [8], based on the previous papers [9, 10, 11], for more details about the argument. We also refer the reader to [17, Chapter 3.5] for a complete review about QED at finite temperature.
@@
In Section 3, the first part of 2.2 stating the well-posedness of the vacuum free energy functional is proved. Then, thanks to additional gauge-invariant estimates, we give a proof of the second part of the theorem providing the main properties of the addressed functional and allowing us to extend it to the most natural function space for this problem. In particular, we find replacements similar to (3) which need to be applied to the null temperature setting [28, 31] in order to get the right calculations for the non-zero temperature case. Due to these prescriptions, the integration by parts with respect to the discretised variable does not hold anymore adding some additional technical difficulties to this problem. Partly for this reason, we restrict our study to the purely magnetic case. Moreover, 3.2 shows that we cannot expect an effect similar to the one shown in [33] for the reduced Bogoliubov-Dirac-Fock (BDF) model, where the authors prove that the thermal effects make the particles rearrange in the polarised Dirac sea in order to completely screen the external potential. This is a classical phenomenon in the field of non-relativistic fermionic plasma physics, known as Debye screening.
Later, in Section 4, we provide the first rigorous derivation of the purely magnetic Euler-Heisenberg formula for the vacuum free energy at positive temperature, in the regime where $B$ is slowly varying in space. In this limit, we prove the convergence of the free energy to the null temperature Euler-Heisenberg formula corrected by some thermal contributions, by recovering a Pauli-Villars regularised effective Lagrangian which coincides with the one first derived by Dittrich [16] in $1979$ (see (4) and (5)), up to some standard mathematical manipulations needed so as to make the integral meaningful as it will be discussed in this section. Even though the null temperature case has been rigorously studied in [31], to the best of our knowledge the non-zero temperature problem has been treated only in the physics literature. More specifically, we consider a magnetic field of the form $B\left(\varepsilon x\right)$ and look at the limit $\varepsilon\xrightarrow{}0$. Notice that in this context we get a strong magnetic potential given by $A_{\varepsilon}\left(x\right)=\varepsilon^{-1}A\left(\varepsilon x\right)$. Nevertheless, this is not a source of additional difficulties from a spectral point of view (see 2.1). Indeed, briefly speaking, the presence of a purely magnetic field ensures us to keep the spectrum of the Dirac operator far apart from zero without adding any additional hypothesis on the size of the potential (as we would expect to do if an electric potential was turned on). In this regime, as stated in 2.3 below, we recover an asymptotic Euler-Heisenberg formula for the vacuum free energy of the following form:
@@ 2. Derivation of the model and main results
(where $C$ is the charge-conjugation operator in Fock space), when the integrand in (6) is interpreted as follows:
@@
$\alpha_{k}=\begin{bmatrix}0&\sigma_{k}\\ \sigma_{k}&0\end{bmatrix},\quad\beta=\begin{bmatrix}I_{2}&0\\ 0&I_{2}\end{bmatrix},$
@@
$\dot{H}^{1}_{\mathrm{div}}\left(\mathbb{R}^{3}\right)=\left\{\boldsymbol{A}=\left(V,A\right)\in L^{6}\left(\mathbb{R}^{3},\mathbb{C}^{4}\right)\mathrel{}\middle|\mathrel{}\mathrm{div}A=0\text{ and }\boldsymbol{F}=\left(-\nabla V,\mathrm{curl}A\right)\in L^{2}\left(\mathbb{R}^{3},\mathbb{R}^{6}\right)\right\},$
@@
The proof of the following lemma, which recalls some spectral properties of the electromagnetic Dirac operator $D^{e\boldsymbol{A}}_{m}$, can be found in [28, Lemma 2.1]:
@@
    The operator $D^{e\boldsymbol{A}}_{m}$ is self-adjoint and its essential spectrum is
@@
The expectation value of the energy in any state in Fock space can be expressed as
$\langle\mathbb{H}^{e\boldsymbol{A}}\rangle=\operatorname{tr}\left[D^{e\boldsymbol{A}}_{m}\left(\gamma-\frac{1}{2}\right)\right]=E\left(\gamma\right),$ (7)
@@
The renormalised density matrix in (7) is a consequence of charge-conjugation invariance. The interested reader can find more details in [35]. Since electrons are fermions, they obey the Pauli exclusion principle, which implies that $\gamma$ has to satisfy $0\leq\gamma\leq 1$ on $L^{2}\left(\mathbb{R}^{3},\mathbb{C}^{4}\right)$. On the other hand, any operator $\gamma$ verifying $0\leq\gamma\leq 1$ arises formally from one state in Fock space. Since the energy depends on the state of the electrons only through the density matrix $\gamma$, we can reformulate our problem by focusing on the simpler operator $\gamma$ and the corresponding energy (7).
@@
$F\left(\gamma\right)=E\left(\gamma\right)+T\operatorname{tr}\left[S\left(\gamma\right)\right],$ (8)
@@
$S\left(x\right)=x\log x+\left(1-x\right)\log\left(1-x\right),$
and the minimisation of the free energy functional with respect to $\gamma$ has to be addressed. For atoms and molecules, we should add a charge constraint of the following form:
@@
In this section, we set the electric potential $V$ to be equal to 00 and focus on the derivation of Dirac’s vacuum free energy in an external magnetic field at positive temperature $T$. Thus, we have not any other constraint than $0\leq\gamma\leq 1$. The reason of this choice will be clearer in the following sections, but essentially working with a magnetic Dirac operator ensures us to keep the spectrum far away from 00 without adding some additional hypothesis on the size of the potential and allowing us to localise the energy to sets of fixed size. By [26, Lemma 1.1], we know that the only critical point $\gamma^{*}$ of functional (8), which is also the unique minimiser by convexity, is given by
@@
$F\left(\gamma^{*}\right)$ $=\operatorname{tr}\left[D^{eA}_{m}\left(\gamma^{*}-\frac{1}{2}\right)\right]+T\operatorname{tr}\left[S\left(\gamma^{*}\right)\right]$
@@
By subtracting an infinite constant, we do not formally change the variational problem that we are interested in, so we also do not change the resulting equations. However, the functional (10) is not well defined yet because of the well known ultraviolet divergence of this theory. Indeed, the operator $\log\left(2\cosh\left(\beta D^{0}_{m}/2\right)\right)-\log\left(2\cosh\left(\beta D^{eA}_{m}/2\right)\right)$ is not trace class as long as $A\neq 0$. This can formally be seen by expanding the trace in a power series of $eA$. As we will see later, the first order term vanishes and the second order term is infinite as long as no ultraviolet cut-off is imposed.
@@
Here, we will use the regularisation technique introduced by Pauli and Villars [45] in $1949$. Of course, this is not the only possible choice when gauge invariance has to be conserved (see for instance [39] for an alternative approach). The Pauli-Villars method consists in introducing $J$ fictitious particles into the model with very high masses $m_{1},\ldots,m_{J}$ playing the role of ultraviolet cut-offs. Indeed, note that $m$ has the dimension of a momentum since in our system of units $\hbar=c=1$. The only role of these additional particles is to regularise the model at high energies and they have not a physical meaning. Indeed, these particles being really massive, they do not affect the low energy regime. In our framework, the Pauli-Villars method consists in introducing the following energy functional:
@@
It is well known in physics literature [45, 32, 5] that only two auxiliary fields are necessary to fulfill these conditions, hence we shall take $J=2$ in the rest of this work. In this case, condition (12) is equivalent to
@@
Of course, once the value of $\Lambda$ is fixed, we cannot uniquely determine $m_{1},m_{2}$. Practically speaking, we usually choose the masses as functions of $\Lambda$ such that the Pauli-Villars coefficients $c_{1},c_{2}$ stay bounded when $\Lambda\xrightarrow{}\infty$. As already mentioned, the logarithmic divergence in the averaged cut-off $\Lambda$ can explicitly be seen in the second order term in the expansion which will be studied in detail in the sequel of this work. A renormalisation of the elementary charge $e$ is necessary in order to remove this remaining divergence, but we will not address this problem here.
Now, for technical reasons, that is to make our computations simpler, we would like to rewrite the Pauli-Villars-regularised free energy functional (11) in an integral form. In order to do so, we derive (11) with respect to $\beta$ by getting:
@@
We can now state one the main results of this paper stating the well-posedness of the PV-regularised free energy functional $\mathcal{F}_{\mathrm{PV}}$ and giving its properties. The proof of this theorem will be given in Section 3.
@@
    For any $A\in L^{1}\left(\mathbb{R}^{3},\mathbb{R}^{3}\right)\cap\dot{H}^{1}_{\mathrm{div}}\left(\mathbb{R}^{3}\right)$, the operator $\operatorname{tr}_{\mathbb{C}^{4}}T_{A}\left(\beta\right)$ is trace class on $L^{2}\left(\mathbb{R}^{3},\mathbb{C}\right)$, well defined with respect to $\beta$ and bounded in a neighbourhood of $\beta=0$. In particular, $\mathcal{F}_{\mathrm{PV}}\left(A,\beta\right)$ is well defined in this case by
    $\mathcal{F}_{\mathrm{PV}}\left(A,\beta\right)=\frac{1}{\beta}\int_{0}^{\beta}\operatorname{tr}\left(\operatorname{tr}_{\mathbb{C}^{4}}T_{A}\left(b\right)\right)\mathop{}\mathopen{}\mathrm{d}b.$
@@
    where $\Lambda$ is defined by (13). Moreover, $M^{T}$ is well defined, bounded and such that
@@
The function $M^{0}$ describes the linear response of the Dirac sea at zero temperature and is very well known in physics literature [32, Eq. (5.39)]. From a mathematical point of view, it has been proved in [28] that
@@
where the function $U$ in the right-hand side was first computed by Serber [49] and Uehling [53]. The same function appears in previous mathematical works studying the purely electrostatic case [36, 34, 30]. As already mentioned, the Pauli-Villars regularisation allows us to remove only the worst linear ultraviolet divergence. Indeed, the second order term is still logarithmically divergent in the high frequencies domain because of $M^{0}$, as shown in (16). On the other hand, $M^{T}$ is the contribution at finite temperature $T$ to the linear response of Dirac’s vacuum. Estimate (17) guarantees $\mathcal{F}_{\mathrm{PV}}$ to be a well defined energy functional from both a mathematical and a physical point of view. For all of these reasons, 2.2 is the equivalent of [28, Theorem 2.1] at positive temperature $T$ and will be discussed much more in detail in Section 3.

### v2

# The Pauli-Villars regularised free energy of Dirac’s vacuum in purely magnetic fields

@@ Abstract.
The Dirac vacuum is a non-linear polarisable medium rather than an empty space. This non-linear behaviour starts to be significant for extremely large electromagnetic fields such as the magnetic field on the surface of certain neutron stars. Even though the null temperature case was deeply studied in the past decades, the problem at non-zero temperature needs to be better understood.
In this work, we present the first rigorous derivation of the one-loop effective magnetic Lagrangian at positive temperature, a non-linear functional describing the free energy of the Dirac vacuum in a classical magnetic field. After introducing our model, we properly define the free energy functional using the Pauli-Villars regularisation technique in order to remove the worst ultraviolet divergences, which represent a well known issue of the theory. The study of the properties of this functional is addressed before focusing on the limit of slowly varying classical magnetic fields. In this regime, we prove the convergence of this functional to the Euler-Heisenberg formula with thermal corrections, recovering the effective Lagrangian first derived by Dittrich [17] in $1979$.
@@ 1. Introduction
The non-linear polarisable nature of the vacuum in quantum field theory has been known for a long time [15, 24, 58]. In this framework, the classical action is replaced by an effective action taking the quantum corrections into account while guaranteeing the validity of the principle of least action. In quantum electrodynamics, a standard way to get the effective action is by integrating out fermions in the path integral (see for instance [50, Ch. 33]). In this manner, we obtain a functional of a classical electromagnetic field treated as an external one. In [30], the authors derive in a rigorous way the effective Lagrangian action for time-independent fields in the Coulomb gauge,
$\mathcal{L}\left(\boldsymbol{A}\right)=-\mathcal{F}_{\mathrm{vac}}\left(e\boldsymbol{A}\right)+e\int_{\mathbb{R}^{3}}\left(j_{\mathrm{ext}}\cdot A\left(x\right)-\rho_{\mathrm{ext}}\left(x\right)V\left(x\right)\right)\mathop{}\mathopen{}\mathrm{d}x\\ +\frac{1}{8\pi}\int_{\mathbb{R}^{3}}\left(|E\left(x\right)|^{2}-|B\left(x\right)|^{2}\right)\mathop{}\mathopen{}\mathrm{d}x,$
@@
where $f_{\mathrm{vac}}\left(eE,eB\right)$ is the energy per unit volume for a constant electromagnetic field $\boldsymbol{F}=\left(E,B\right)$. The function $f_{\mathrm{vac}}$ was first computed by Euler and Heisenberg [40] in $1936$. Some alternative derivations were then presented by Weisskopf [58] and Schwinger [51], who introduced a largely renowned method in the physics literature known as proper time method. Since the integrand in the Euler-Heisenberg formula [34, Formula 4] can have poles on the real line, a proper definition requires to shift the integration path to the complex plane by replacing the integral variable $s$ by $s+i\eta$ and taking the limit $\eta\xrightarrow{}0$ [51]. In this manner, the function $f_{\mathrm{vac}}$ gains an exponentially small non-zero imaginary part which was interpreted by Schwinger as the electron-positron pair production rate and explains the instability of the vacuum [18][35, Paragraph 7.3]. Nevertheless, in the case of a purely magnetic field, the Euler-Heisenberg formula is absolutely convergent and real, as expected. For this reason, this situation has been deeply studied in the physics literature [8, 54, 55, 44] and the function $f_{\mathrm{vac}}$ can be simplified as follows:
@@
In the non-perturbative physical examples above, the fields exist in a thermal bath or in some non-equilibrium background which is very different from the vacuum. For this reason, the study of finite temperature effects in quantum field theory attracted the interest of many physicists in the past decades. In particular, several authors [41, 57, 19] considered what happens when a system of elementary particles described by a quantum field theory is heated. They found that symmetries which are spontaneously broken at zero temperature (such as those of the weak interactions) may be restored at sufficiently high temperatures, and calculated the critical temperature at which such a restoration takes place. To do this kind of calculation, one needs to know the Feynman rules for a field theory at finite temperature. For a non-gauge theory, these rules can be derived using well known methods [26]. However, for a gauge theory, a more powerful technique is needed to cope with several new problems that arise [5]. In any case, the Feynman rules at finite temperature $T$ are strictly related to the Kubo-Martin-Schwinger (KMS) condition (see 3.4) and consist in the zero temperature rules with the following formal replacements:
@@
It is now clear that the calculation of the effective action at positive temperature has been an interesting physical question for some decades, particularly because the imaginary part of the effective action is expected to give the vacuum decay rate and one would like to understand the influence of the temperature on this rate. There have been several attempts to generalise Schwinger’s proper time method to the finite temperature case in order to determine the temperature dependent effective actions, leading to conflicting results in the physics literature. The subtleties behind these results have to do with the imposition of the anti-periodicity condition required at positive temperature as well as on the finite temperature formalism used to carry out the calculation. We refer the interested reader to [9], based on the previous papers [10, 11, 12], for more details about the argument. We also refer the reader to [18, Chapter 3.5] for a complete review about QED at finite temperature.
@@
In Section 3, the first part of 2.2 stating the well-posedness of the vacuum free energy functional is proved. Then, thanks to additional gauge-invariant estimates, we give a proof of the second part of the theorem providing the main properties of the addressed functional and allowing us to extend it to the most natural function space for this problem. In particular, we find replacements similar to (3) which need to be applied to the null temperature setting [30, 34] in order to get the right calculations for the non-zero temperature case. Due to these prescriptions, the integration by parts with respect to the discretised variable does not hold anymore adding some additional technical difficulties to this problem. Partly for this reason, we restrict our study to the purely magnetic case. Moreover, 3.3 shows that we cannot expect an effect similar to the one shown in [36] for the reduced Bogoliubov-Dirac-Fock (BDF) model, where the authors prove that the thermal effects make the particles rearrange in the polarised Dirac sea in order to completely screen the external potential. This is a classical phenomenon in the field of non-relativistic fermionic plasma physics, known as Debye screening.
Later, in Section 4, we provide the first rigorous derivation of the purely magnetic Euler-Heisenberg formula for the vacuum free energy at positive temperature, in the regime where $B$ is slowly varying in space. In this limit, we prove the convergence of the free energy to the null temperature Euler-Heisenberg formula corrected by thermal contributions, recovering a Pauli-Villars regularised effective Lagrangian which coincides with the one first derived by Dittrich [17] in $1979$ (see (4) and (5)), up to some standard mathematical manipulations needed so as to give a rigorous meaning to this object, as it will be discussed in this section. Even though the null temperature case has been rigorously studied in [34], to the best of our knowledge the non-zero temperature problem has been treated only in the physics literature. More specifically, we consider a magnetic field of the form $B\left(\varepsilon x\right)$ and look at the limit $\varepsilon\xrightarrow{}0$. Notice that in this context we get a strong magnetic potential given by $A_{\varepsilon}\left(x\right)=\varepsilon^{-1}A\left(\varepsilon x\right)$. Nevertheless, this is not a source of additional difficulties from a spectral point of view (see 2.1). Indeed, briefly speaking, the presence of a purely magnetic field ensures us to keep the spectrum of the Dirac operator far apart from zero without adding any additional hypothesis on the size of the potential (as we would expect to do if an electric potential was turned on). In this regime, as stated in 2.3 below, we recover an asymptotic Euler-Heisenberg formula for the vacuum free energy of the following form:
@@ 2. Derivation of the model and main results
(where $C$ is the charge-conjugation operator in Fock space as defined in [53, Formula (1.81)]), when the integrand in (6) is interpreted as follows:
@@
$\alpha_{k}=\begin{bmatrix}0&\sigma_{k}\\ \sigma_{k}&0\end{bmatrix},\quad\beta=\begin{bmatrix}I_{2}&0\\ 0&-I_{2}\end{bmatrix},$
@@
$\dot{H}^{1}_{\mathrm{div}}\left(\mathbb{R}^{3}\right)=\left\{\boldsymbol{A}=\left(V,A\right)\in L^{6}\left(\mathbb{R}^{3},\mathbb{R}^{4}\right)\mathrel{}\middle|\mathrel{}\mathrm{div}A=0\text{ and }\boldsymbol{F}=\left(-\nabla V,\mathrm{curl}A\right)\in L^{2}\left(\mathbb{R}^{3},\mathbb{R}^{6}\right)\right\},$
@@
The proof of the following lemma, which recalls some spectral properties of the electromagnetic Dirac operator $D^{e\boldsymbol{A}}_{m}$, can be found in [30, Lemma 2.1]:
@@
    The operator $D^{e\boldsymbol{A}}_{m}$ is self-adjoint on $H^{1}\left(\mathbb{R}^{3},\mathbb{C}^{4}\right)$ and its essential spectrum is
@@
The expectation value of the energy in any state in Fock space can be expressed as [30, Equation (2.8)]
$\langle\mathbb{H}^{e\boldsymbol{A}}\rangle=\operatorname{tr}_{L^{2}\left(\mathbb{R}^{3},\mathbb{C}^{4}\right)}\left[D^{e\boldsymbol{A}}_{m}\left(\gamma-\frac{1}{2}\right)\right]=E\left(\gamma\right),$ (7)
@@
The renormalised density matrix in (7) is a consequence of charge-conjugation invariance. The interested reader can find more details in [38]. Since electrons are fermions, they obey the Pauli exclusion principle, which implies that $\gamma$ has to satisfy $0\leq\gamma\leq 1$ on $L^{2}\left(\mathbb{R}^{3},\mathbb{C}^{4}\right)$. On the other hand, any operator $\gamma$ verifying $0\leq\gamma\leq 1$ formally arises from one state in Fock space. Since the energy depends on the state of the electrons only through the density matrix $\gamma$, we can reformulate our problem by focusing on the simpler operator $\gamma$ and the corresponding energy (7).
@@
$F\left(\gamma\right)=E\left(\gamma\right)-T\operatorname{tr}\left[S\left(\gamma\right)\right],$ (8)
@@
$S\left(x\right)=-x\log x-\left(1-x\right)\log\left(1-x\right)$
and therefore $\operatorname{tr}\left[S\left(\gamma\right)\right]$ is an entropy term which is well known in the literature (see e.g. [2, Formula (2c.10)], [36, Formula (4)] and [28, Formula (5)]). Then, the minimisation of the free energy functional with respect to $\gamma$ has to be addressed. For atoms and molecules, one should add a charge constraint of the following form:
@@
Nevertheless, in this work, we restrict ourselves to the vacuum case for simplicity. Therefore, we do not have any other constraint than $0\leq\gamma\leq 1$. In this section, we set the electric potential $V$ to be equal to 00 and focus on the derivation of Dirac’s vacuum free energy in an external magnetic field at positive temperature $T$. The reason of this choice will be clearer in the following sections, but essentially working with a magnetic Dirac operator ensures us to keep the spectrum far away from 00 without adding some additional hypothesis on the size of the potential and allowing us to localise the energy to sets of fixed size. By [28, Lemma 1.1], we know that the only critical point $\gamma^{*}$ of functional (8), which is also the unique minimiser by convexity, is given by
@@
$F\left(\gamma^{*}\right)$ $=\operatorname{tr}\left[D^{eA}_{m}\left(\gamma^{*}-\frac{1}{2}\right)\right]-T\operatorname{tr}\left[S\left(\gamma^{*}\right)\right]$
@@
By subtracting an infinite constant, we do not formally change the variational problem that we are interested in, so we also do not change the resulting equations. However, the functional (10) is not well defined yet because of the well known ultraviolet divergences of this theory. Indeed, the operator $\log\left(2\cosh\left(\beta D^{0}_{m}/2\right)\right)-\log\left(2\cosh\left(\beta D^{eA}_{m}/2\right)\right)$ is not trace class as long as $A\neq 0$. This can formally be seen by expanding the trace in a power series of $eA$. As we will see later, the first order term vanishes and the second order term is infinite due to ultraviolet divergences. Moreover, the formal higher order terms cannot prevent this ultraviolet divergence.
@@
Here, we will use the regularisation technique introduced by Pauli and Villars [48] in $1949$. Of course, this is not the only possible choice when gauge invariance has to be conserved (see for instance [42] for an alternative approach). The Pauli-Villars method consists in introducing $J$ fictitious particles into the model with very high masses $m_{1},\ldots,m_{J}$ playing the role of ultraviolet cut-offs. Indeed, note that $m$ has the dimension of a momentum since in our system of units $\hbar=c=1$. The only role of these additional particles is to regularise the model at high energies and they have no physical meaning. Indeed, these particles being really massive, they do not affect the low energy regime. In our framework, the Pauli-Villars method consists in introducing the following energy functional:
@@
It is well known in the physics literature [48, 35, 6] that only two auxiliary fields are necessary to fulfill these conditions, hence we shall take $J=2$ in the rest of this work. In this case, condition (12) is equivalent to
@@
Of course, once the value of $\Lambda$ is fixed, we cannot uniquely determine $m_{1},m_{2}$. Practically speaking, we usually choose the masses as functions of $\Lambda$ such that the Pauli-Villars coefficients $c_{1},c_{2}$ stay bounded when $\Lambda\xrightarrow{}\infty$. As already mentioned, the logarithmic divergence in the averaged cut-off $\Lambda$ can explicitly be seen in the second order term in the expansion which will be studied in detail in the sequel of this work (see Section 3.2). A renormalisation of the elementary charge $e$ is necessary in order to remove this remaining divergence, but we will not address this problem here.
Now, in order to simplify our calculations, we rewrite the Pauli-Villars-regularised free energy functional in an integral form. In order to do so, we derive (11) with respect to $\beta$ by getting:
@@
We can now state one of the main results of this paper stating the well-posedness of the PV-regularised free energy functional $\mathcal{F}_{\mathrm{PV}}$ and giving its properties. The proof of this theorem will be given in Section 3.
@@
    For any $A\in L^{1}\left(\mathbb{R}^{3},\mathbb{R}^{3}\right)\cap H^{1}\left(\mathbb{R}^{3},\mathbb{R}^{3}\right)$, the operator $T_{A}\left(\beta\right)$ is trace class on $L^{2}\left(\mathbb{R}^{3},\mathbb{C}^{4}\right)$, well defined with respect to $\beta$ and bounded in a neighbourhood of $\beta=0$. In particular, $\mathcal{F}_{\mathrm{PV}}\left(A,\beta\right)$ is well defined in this case by
    $\mathcal{F}_{\mathrm{PV}}\left(A,\beta\right)=\frac{1}{\beta}\int_{0}^{\beta}\operatorname{tr}\left[T_{A}\left(b\right)\right]\mathop{}\mathopen{}\mathrm{d}b.$
@@
    where $\Lambda$ is defined by (13). Moreover, $M^{T}$ is well defined, bounded for any fixed $\beta\in\mathopen{(}0\mathclose{}\mathpunct{},+\infty\mathclose{)}$ and such that
@@
The function $M^{0}$ describes the linear response of the Dirac sea at zero temperature and is very well known in the physics literature [35, Eq. (5.39)]. From a mathematical point of view, it has been proved in [30] that
@@
where the function $U$ in the right-hand side was first computed by Serber [52] and Uehling [56]. The same function appears in previous mathematical works studying the purely electrostatic case [39, 37, 33]. As already mentioned, the Pauli-Villars regularisation allows us to remove only the worst linear ultraviolet divergences. Indeed, the second order term is still logarithmically divergent in the high frequencies domain because of $M^{0}$, as shown in (16). On the other hand, $M^{T}$ is the contribution at finite temperature $T$ to the linear response of Dirac’s vacuum. Estimate (17) guarantees $\mathcal{F}_{\mathrm{PV}}$ to be a well defined free energy functional from both a mathematical and a physical point of view. For all of these reasons, 2.2 is the equivalent of [30, Theorem 2.1] at positive temperature $T$ and will be discussed much more in detail in Section 3.